OpenStudy (anonymous):

Please help I don't understand this at all!! Find the length of the hypotenuse of an isosceles right triangle whose legs are 1 unit in length.

OpenStudy (here_to_help15):

a^2+b^2=c^2 1^2+1^2=2 c=sq root of 2=1.414